OpenStudy (anonymous):

simplify 49 ^-1/2 1/7 -1/7 7 -7

OpenStudy (angela210793):

=1/49^1/2=1/sqr49=1/7 but can also b -1/7

OpenStudy (anonymous):

so which one do i choose?

OpenStudy (anonymous):

49^(-1/2) = 1/sqrt{49} = 1/7

OpenStudy (angela210793):

hmmmmm 1/7 i'd say....u see 4 example sqr4=+-2

OpenStudy (anonymous):

yes